# ImplicitQA: Going beyond frames towards Implicit Video Reasoning

### v1

@@ Abstract
[…]
Figure 1: ImplicitQA examples, each targeting a distinct implicit-reasoning dimension. (a) Lateral spatial reasoning—identifying the toy opposite the wizard clock by mentally mapping objects across the scene. (b) Motion and trajectory dynamics—inferring that black bullets move away from Mario by integrating actions and character positions. (c) Inferred counting—determining which animal is the third to leave a bridge by tracking sequential departures that are never fully visible onscreen. Models that excel at explicit perception often fail on these tasks, highlighting the need for benchmarks that probe deeper narrative understanding.
@@ 1 Introduction
Video Question Answering (VideoQA) sits at the intersection of computer vision and natural language processing, aiming to answer natural language questions based on visual content in videos. Recent progress in VideoQA [13, 20, 9, 15, 5, 6, 21, 14, 1, 19] has been fueled by multimodal learning techniques that integrate visual and textual modalities, enabling impressive performance on datasets where questions are grounded in explicit visual content. These benchmarks [17, 18, 16, 4, 7, 2, 3, 12, 8] typically emphasize recognizing objects, identifying actions, and understanding events that are directly observable within individual frames or short clips. However, cinematic and narrative-driven videos - such as movies frequently employ storytelling devices that challenge this paradigm. Rather than explicitly showing every key detail, such videos rely on indirect cues, subtle scene transitions, and off-screen implications to advance the plot or convey meaning. For example, a character’s motive might be implied through a prior conversation but never visually depicted, or a causal event might occur off-screen, requiring viewers to infer its effects from subsequent scenes. Humans are remarkably skilled at making such implicit inferences, seamlessly connecting context, prior knowledge, and temporal cues to reconstruct missing narrative elements. This ability forms a cornerstone of real-world video understanding. However, current VideoQA benchmarks largely overlook this aspect of human-like reasoning. They reward models that excel at surface-level perception but do not challenge systems to infer beyond what is visually present. As a result, today’s state-of-the-art models may perform well on existing datasets while failing to grasp deeper, implicit reasoning. Consider the example depicted in Figure 1(b), where the question posed is: "In what direction were the black bullets heading relative to the man in the red hat?" with answer choices: {’A’: ’Towards him’, ’B’: ’Away from him’, ’C’: ’To his right’, ’D’: ’To his left’}. The correct answer is "B" (Away from him), as the scene implicitly shows the princess Peach running towards Mario while black bullets are fired towards her, thus moving away from Mario. However, existing VideoQA models consistently answer incorrectly, choosing "Towards him," failing to grasp the information across frames. Even the O3 model incorrectly selects "To his left," highlighting the widespread difficulty in capturing implicit narrative dynamics.
@@
To address this critical gap, we introduce ImplicitQA, a benchmark designed to probe the limits of implicit reasoning in VideoQA. We collect $1K$ carefully curated question-answer pairs from 320+ diverse video clips sourced from movies. Unlike traditional benchmarks, ImplicitQA focuses exclusively on questions that cannot be answered through direct observation of frames alone. Instead, they require reasoning about unstated causality, character motives, social interactions, physical context, and other narrative nuances that are implied but not explicitly depicted.
@@
Our dataset is organized into $9$ core reasoning categories: (1) Lateral Spatial Reasoning, (2) Vertical Spatial Reasoning, (3) Relative Depth and Proximity, (4) Motion and Trajectory Dynamics, (5) Viewpoint and Visibility, (6) Causal and Motivational Reasoning, (7) Social Interaction and Relationships, (8) Physical and Environmental Context, and (9) Inferred Counting. We have annotated the data ourselves, thus ensuring that the questions are both challenging and aligned with the nuanced reasoning capabilities we aim to benchmark. Further, we re-verify the annotations amongst ourselves, this curation process also minimizes the ambiguity and guarantees that each question tests meaningful aspects of implicit understanding.
@@
We evaluate on prominent VideoQA models on ImplicitQA and observe significant performance drops compared to standard benchmarks. This finding underscores the current limitations of VideoQA systems, which remain heavily reliant on explicit visual cues. Notably, we find that reasoning-oriented models outperform non-reasoning models: for example, GPT-o3[11] achieves a 9.8% higher accuracy than GPT-4.1[10]. This gap illustrates the necessity of deeper reasoning capabilities to tackle the challenges posed by ImplicitQA, further validating the benchmark’s focus.
@@
  • We manually curate a high-quality dataset of 1k question-answer pairs across 320+ diverse video clips, with annotation conducted by experts in computer vision to ensure rigor and relevance.
  • We define a taxonomy of 9 categories, covering lateral spatial reasoning, depth and proximity, causal inference, social dynamics, and more, to facilitate targeted analysis and benchmarking.
@@
  • We release both the dataset and the data collection framework, providing tools for reproducibility and encouraging further research.
@@ 2 Related Benchmarks
We compare our proposed datasets against some recent VideoQA works, which we broadly categorize into Vision-only benchmarks, and Vision and speech benchmarks, which require integration of information from othe modalities like speech.
@@
Cinepile[12] Movie Clips QA (Templates + LLM) LLM
TVQA[4] TV Show Clips QA (Templates) - -
VideoMME[3] Diverse Experts - LLM
MVBench[7] Prior Datasets Prior Datasets Wrong Options -
TemporalBench[2] Prior Datasets Captions Pairing LLM
TempCompass[8] Stock + Transforms Class Labels QA (Templates + LLM) -
@@ 2.1 Vision Only Benchmarks
MVBench [7]: MVBench aggregates roughly 4,000 human- or automatically–derived multiple-choice questions drawn from 11 public video and image datasets, pairing each 5–35 s clip with four or five answer options. A “static-to-dynamic” pipeline converts image-based tasks into temporally grounded ones, ensuring coverage of both short-term motions and longer-horizon causal phenomena. Prompts are minimal—models output only the option letter for deterministic scoring. While MVBench systematically probes temporal understanding, many items still admit answers via single-frame cues or rely on source annotations not originally crafted for VideoQA.
TempCompass [8]: TempCompass focuses on pure temporal manipulation by algorithmically editing 410 royalty-free clips ($\leq$30 s) into conflictual pairs/triplets that differ only in one temporal property (e.g., playback reversal, reordered events, speed changes). GPT-3.5 then generates 7,540 diverse tasks (multiple-choice, yes/no, caption matching, constrained captioning) from concise “meta cards.” Although it powerfully discourages single-frame shortcuts, its short clips limit long-range reasoning, and the automated grading pipeline (rule checks + LLM evaluator) can mis-score free-form outputs.
TemporalBench [2]: TemporalBench curates 2,032 silent clips (<20 s) from seven public corpora, enriches each with a dense human-written caption plus up to 15 machine-crafted counterfactual captions, and forms 9,867 contrastive pairs targeting fine-grained temporal distinctions (action order, frequency, direction, effector). Its dual “short” vs. “long” settings (up to 20 min concatenations) enable both local and extended temporal tests. By framing every item as a binary choice, it blocks majority-vote biases—but the absence of linguistic context and reliance on caption edits still leave room for non-visual shortcuts.
ImplicitQA is designed to go beyond temporal or multimodal video benchmarks: Firstly, Implicit, multi-frame inference. Questions demand reasoning about off-screen events, unstated motives, or causal chains spanning multiple clips—no single-frame or caption-only cues suffice. Secondly, we utilize expert manual curation. Unlike MVBench’s mixed annotations, TempCompass’s algorithmic edits, or TemporalBench’s counterfactual pipeline, our 1K questions over 320+ clips are manually authored and rigorously cross-verified by computer-vision experts, ensuring each item truly probes implicit visual reasoning.
@@ 2.2 Vision and Speech Fusion Benchmarks
VideoMME [3] VideoMME comprises 900 expert-annotated YouTube videos (short <2 min, medium 4–15 min, long 30–60 min) with frames, raw audio, and automatically extracted subtitles. Human annotators author balanced multiple-choice questions across diverse domains, and a “text-only” filter removes any item solvable without visual or acoustic cues—though many still lean on subtitle or audio information.
TVQA [4] TVQA provides $\sim$ 22 K clips (60–90 s) from six U.S. TV series, each paired with dialogue subtitles and precise timestamps. Its compositional “WH-word … when/before/after …” templates force joint vision–language reasoning and moment localization, but the narrow domain and reliance on spoken dialogue limit purely visual inference.
CinePile [12] CinePile curates $\sim$ 9K movie snippets ($\sim$ 160 s) from MovieClips, augmented with professional audio descriptions and subtitles. An LLM-driven pipeline yields 304 K multiple-choice QAs, filtered adversarially for shortcut resistance. Despite its narrative breadth, the fully automated QA creation can introduce noise and remains tied to explicit descriptions.
ImplicitQA departs from these benchmarks in three key ways: Firstly, implicit multi-frame inference: every question demands reasoning about off-screen events, unstated motives, or causal chains—not solvable by a single frame or subtitle snippet. Secondly, Visual-only focus: We strip away subtitles and audio tracks entirely—models must extract and integrate visual cues across frames without any textual crutch. Finally, expert manual curation: Unlike datasets that rely on automated pipelines or narrow scripted dialogue, our 1 K questions across 320+ clips are manually authored and cross-verified by computer-vision experts, ensuring each item truly probes implicit visual reasoning.
@@ 3.1 Annotation Tool and Data Collection
Figure 2: ImplicitQA Curation Pipeline. We begin by selecting creative video clips and download them. An expert-annotator pool then uses our FrameQuiz Annotation Tool to (1) mark temporal segments, (2) add a multiple-choice question and its correct answer for the segment, and (3) craft plausible distractor options. These annotated clips form the raw ImplicitQA Dataset. Next, a non-expert annotator pool employs the ImplicitEval Annotation Tool to answer each question, yielding a human baseline accuracy score. We run GPT-4.1[10] on the dataset to automatically assign initial category tags, which are then relabeled by the expert annotators.
To ensure annotation quality and reliability, we implemented a save-and-replay feature within the tool, enabling annotators to revisit their annotated segments, replay the selected video portion, and iteratively refine or validate their annotations before final submission. We also verify the annotations amongst ourselves. This process ensured that questions accurately aligned with the video context and targeted implicit rather than explicit reasoning. We ourselves have annotated questions with the intent to probe deeper inferential reasoning rather than relying solely on directly observable content. Thus the author annotation process contributed to both the conceptual depth and technical relevance of the dataset.For video selection, we curated a diverse set of 320+ creative videos, comprising movies of varied genres and mediums (3D animated and live-action) known for employing narrative techniques such as implied causality, off-screen action, symbolic representation, and indirect storytelling. We prioritized scenes that challenge viewers to make inferences beyond directly visible actions or objects, focusing on content where critical narrative elements are omitted, subtle, or distributed across frames. Statistical characteristics of our collected dataset can be seen in Figure 3.
@@ 3.2 Dataset Categorization
To ensure comprehensive coverage of the diverse reasoning abilities required for implicit video understanding, we organized the dataset into nine distinct reasoning categories, each targeting a specific type of implicit inference. Below, we formally define each category of implicit reasoning. We present qualitative examples in Figure 16 to highlight the implicit nature of questions. We also show differences in temporal length between categories in Figure 4, with counting and vertical spatial reasoning videos being the longest.
@@
Viewpoint and Visibility Questions in this category require inferring what is observable from a particular vantage point - whether it be a character’s perspective or a camera angle. Solvers must reason about line-of-sight, occlusions, and spatial orientations.
@@
Causal and Motivational Reasoning Questions in this category require viewers to infer causal relationships, character motives, or likely future events based on incomplete or indirect visual information. These questions emphasize unstated cause-effect chains within the story.
@@
Figure 6: More Qualitative ImplicitQA examples, targeting a distinct implicit-reasoning dimension.

### v2

@@ Abstract
Video Question Answering (VideoQA) has made significant strides by leveraging multimodal learning to align visual and textual modalities. However, current benchmarks overwhelmingly focus on questions answerable through explicit visual content - actions, objects, and events directly observable within individual frames or short clips. In contrast, creative and cinematic videos - such as movies, TV shows, and narrative-driven content - employ storytelling techniques that deliberately omit certain depictions, requiring viewers to infer motives, relationships across discontinuous frames with disjoint visual contexts. Humans naturally excel at such implicit reasoning, seamlessly integrating information across time and context to construct coherent narratives. Yet current benchmarks fail to capture this essential dimension of human-like understanding. To bridge this gap, we present ImplicitQA, a novel benchmark specifically designed to test VideoQA models on human-like implicit reasoning. ImplicitQA comprises $1K$ meticulously annotated QA pairs drawn from $1K$ high-quality creative video clips covering $15$ genres across $7$ decades of content. Questions are systematically categorized into nine key reasoning dimensions: lateral and vertical spatial reasoning, depth and proximity, viewpoint and visibility, motion and trajectory, causal and motivational reasoning, social interactions, physical context, and inferred counting. These annotations are deliberately challenging, crafted by authors, validated through multiple annotators, and benchmarked against human performance to ensure high quality. Our extensive evaluations on $11$ leading VideoQA models reveals consistent and significant performance degradation, underscoring their reliance on surface-level visual cues and highlighting the difficulty of implicit reasoning. Even the best model substantially underperforms human baselines with only 64% accuracy, and no open-source model exceeds 50% accuracy. Performance variations across models further illustrate the complexity and diversity of the challenges presented by ImplicitQA. Our analysis highlights the unique challenges of implicit reasoning, including limited gains from scaling frames or parameters. By releasing both the dataset and data collection framework, ImplicitQA establishes a rigorous, diverse, and reproducible testbed for advancing VideoQA.
Figure 1: ImplicitQA examples, each targeting a distinct implicit‐reasoning dimension. (a) Lateral spatial reasoning-identifying the toy opposite the wizard clock by mentally mapping objects across the scene. (b) Motion and trajectory dynamics-inferring that black bullets move away from Mario by integrating actions and character positions. (c) Inferred counting-determining which animal is the third to leave a bridge by tracking sequential departures that are never fully visible onscreen. Models that excel at explicit perception often fail on these tasks, highlighting the need for benchmarks that probe deeper narrative understanding.
@@ 1 Introduction
Video Question Answering (VideoQA) sits at the intersection of computer vision and natural language processing, aiming to answer natural language questions based on visual content in videos. Recent progress in VideoQA Ren et al. [2023], Zhang et al. [2024a], Maaz et al. [2024], Wang et al. [2024a], Li et al. [2024a; b], Zhang et al. [2024b], Wang et al. [2024b], Bai et al. [2025], Yuan et al. [2025] has been fueled by multimodal learning techniques that integrate visual and textual modalities, enabling impressive performance on datasets where questions are grounded in explicit visual content. These benchmarks Xu et al. [2016], Yu et al. [2019], Xiao et al. [2021], Lei et al. [2018], Li et al. [2023], Cai et al. [2024], Fu et al. [2024], Rawal et al. [2024], Liu et al. [2024], Swetha et al. [2025] typically emphasize recognizing objects, identifying actions, and understanding events that are directly observable within individual frames or short clips. However, cinematic and narrative-driven videos - such as movies frequently employ storytelling devices that challenge this paradigm. Rather than explicitly showing every key detail, such videos rely on indirect cues, subtle scene transitions, and off-screen implications to advance the plot or convey meaning. For example, a character’s motive might be implied through a prior conversation but never visually depicted, or a causal event might occur off-screen, requiring viewers to infer its effects from subsequent scenes. Humans are remarkably skilled at making such implicit inferences, seamlessly connecting context, prior knowledge, and temporal cues to reconstruct missing narrative elements. This ability forms a cornerstone of real-world video understanding. However, current VideoQA benchmarks largely overlook this aspect of human-like reasoning. They reward models that excel at surface-level perception but do not challenge systems to infer beyond what is visually present. As a result, today’s state-of-the-art models may perform well on existing datasets while failing to grasp deeper, implicit reasoning. Consider the example depicted in Figure 1(b), where the question posed is: "In what direction were the black bullets heading relative to the man in the red hat?" with answer choices: {’A’: ’Towards him’, ’B’: ’Away from him’, ’C’: ’To his right’, ’D’: ’To his left’}. The correct answer is "B" (Away from him), as the scene implicitly shows the princess Peach running towards Mario while black bullets are fired towards her, thus moving away from Mario. However, existing VideoQA models consistently answer incorrectly, choosing "Towards him," failing to grasp the information across frames. Even the O3 model incorrectly selects "To his left," highlighting the widespread difficulty in capturing implicit narrative dynamics.
@@
To address this critical gap, we introduce ImplicitQA, a benchmark designed to probe the limits of implicit reasoning in VideoQA. We collect $1K$ carefully curated question-answer pairs from $1K$ diverse video clips sourced from movies. Unlike traditional benchmarks, ImplicitQA focuses exclusively on questions that cannot be answered through direct observation of frames alone. Instead, they require reasoning about unstated character motives, social interactions, physical context, and other narrative nuances that are implied but not explicitly depicted.
@@
Our dataset is organized into $9$ core reasoning categories: (1) Lateral Spatial Reasoning, (2) Vertical Spatial Reasoning, (3) Relative Depth and Proximity, (4) Motion and Trajectory Dynamics, (5) Viewpoint and Visibility, (6) Motivational Reasoning, (7) Social Interaction and Relationships, (8) Physical and Environmental Context, and (9) Inferred Counting. We have annotated the data ourselves, thus ensuring that the questions are both challenging and aligned with the nuanced reasoning capabilities we aim to benchmark. Further, we re-verify the annotations amongst ourselves, this curation process also minimizes the ambiguity and guarantees that each question tests meaningful aspects of implicit understanding.
@@
We evaluate on prominent VideoQA models on ImplicitQA and observe significant performance drops compared to standard benchmarks. This finding underscores the current limitations of VideoQA systems, which remain heavily reliant on explicit visual cues. Notably, we find that reasoning-oriented models outperform non-reasoning models: for example, GPT-o3OpenAI [2024a] achieves a 9.8% higher accuracy than GPT-4.1OpenAI [2024b]. This gap illustrates the necessity of deeper reasoning capabilities to tackle the challenges posed by ImplicitQA, further validating our focus.
@@
  • We manually curate a high-quality dataset of 1k QA pairs across 1k diverse video clips, with annotation conducted by experts in computer vision to ensure rigor and relevance.
  • We define a taxonomy of 9 categories, covering lateral spatial reasoning, depth and proximity, social dynamics, and more, to facilitate targeted analysis and benchmarking.
@@ 2 Related Benchmarks
We compare our proposed dataset against some recent VideoQA benchmarks, which we broadly categorize into Vision-only benchmarks, and Vision and speech benchmarks, which require integration of information from othe modalities like speech.
@@
Cinepile Rawal et al. [2024] Movie Clips QA (Templates + LLM) LLM
TVQA Lei et al. [2018] TV Show Clips QA (Templates) - -
VideoMME Fu et al. [2024] Diverse Experts - LLM
MVBench Li et al. [2023] Prior Datasets Prior Datasets Wrong Options -
TemporalBench Cai et al. [2024] Prior Datasets Captions Pairing LLM
TempCompass Liu et al. [2024] Stock + Transforms Class Labels QA (Templates + LLM) -
Open-EQA Majumdar et al. [2024] Indoor Egocentric tours Human - -
VSI-Bench Yang et al. [2025] ! Indoor Egocentric tours QA (Templates) QA (Templates) Human
@@ 2.1 Vision Only Benchmarks
MVBench Li et al. [2023]: MVBench aggregates roughly 4,000 human- or automatically–derived multiple-choice questions drawn from 11 public age datasets, pairing each 5–35 s clip with four or five answer options. A “static-to-dynamic” pipeline converts image-based tasks into temporally grounded ones, ensuring coverage of both short-term motions and longer-horizon causal phenomena.
TempCompass Liu et al. [2024]: TempCompass focuses on pure temporal manipulation by algorithmically editing 410 royalty-free clips ($\leq$30 s) into pairs/triplets that differ only in one temporal property (e.g., playback reversal, reordered events, speed changes). GPT-3.5 then generates 7,540 diverse tasks (multiple-choice, yes/no, caption matching, constrained captioning).
TemporalBench Cai et al. [2024]: TemporalBench curates 2,032 silent clips (<20 s) from seven public corpora, enriches each with a dense human-written caption plus up to 15 machine-crafted counterfactual captions, and forms 9,867 contrastive pairs targeting fine-grained temporal distinctions (action order, frequency, direction, effector).
Open-EQA Majumdar et al. [2024] and VSIBench Yang et al. [2025]: Both of these datasets are based one pre-existing 3d scans of indoor environments taken from an egocentric POV. As a result the videos in these datasets carry a dense view of the environment, which needs to be stitched together to answer questions. Due to the dense nature of the sampling, these videos primarily require limited implicit spatial reasoning (indicated by !  symbol). VSI-Bench spans configuration, measurement, and spatiotemporal tasks (e.g., room size, appearance order); OpenEQA covers seven categories (object/attribute/state recognition, localization, spatial and functional reasoning, world knowledge).
ImplicitQA is designed to go beyond existing video benchmarks: Firstly, Implicit, multi-frame inference. Questions demand reasoning about off-screen events, unstated motives, or causal chains spanning multiple clips, no single-frame cues suffice. Secondly, we utilize expert manual curation. Unlike MVBench’s mixed annotations, TempCompass’s algorithmic edits, or TemporalBench’s counterfactual pipeline, our 1K questions from 1k clips are manually authored and rigorously cross-verified by computer-vision experts, ensuring each item truly probes implicit visual reasoning.
@@ 2.2 Vision and Speech Fusion Benchmarks
VideoMME Fu et al. [2024] VideoMME comprises 900 expert‐annotated YouTube videos (short <2 min, medium 4–15 min, long 30–60 min) with frames, raw audio, and automatically extracted subtitles. Human annotators author balanced multiple‐choice questions across diverse domains, and a “text‐only” filter removes any item solvable without visual or acoustic cues.
TVQA Lei et al. [2018] TVQA provides $\sim$ 22 K clips (60–90 s) from six U.S. TV series, each paired with dialogue subtitles and precise timestamps. Its compositional “WH‐word … when/before/after …” templates force joint vision–language reasoning and moment localization, but the narrow domain and reliance on spoken dialogue limit purely visual inference.
CinePile Rawal et al. [2024] CinePile curates $\sim$ 9K movie snippets ($\sim$ 160 s) from MovieClips, augmented with professional audio descriptions and subtitles. An LLM‐driven pipeline yields 304 K multiple‐choice QAs, filtered adversarially for shortcut resistance.
ImplicitQA departs from these benchmarks in three key ways: Firstly, implicit multi‐frame inference: every question demands reasoning about off-screen events, unstated motives, or causal chains not solvable by a single frame or subtitle snippet. Secondly, Visual-only focus: We strip away subtitles and audio tracks entirely, models must extract and integrate visual cues across frames without any textual crutch. Finally, expert manual curation: Unlike datasets that rely on automated pipelines or narrow scripted dialogue, our 1K questions across 1K clips are manually authored and cross-verified by computer‐vision experts, ensuring each item truly probes implicit visual reasoning.
@@ 3.1 Annotation Tool and Data Collection
Figure 2: ImplicitQA Curation Pipeline. We begin by selecting creative video clips and download them. An expert‐annotator pool then uses our FrameQuiz Annotation Tool to (1) mark temporal segments, (2) add a multiple‑choice question and its correct answer for the segment, and (3) craft plausible distractor options. These annotated clips form the raw ImplicitQA Dataset. Next, a non‑expert annotator pool employs the ImplicitEval Annotation Tool to answer each question, yielding a human baseline accuracy score. We run GPT‑4.1OpenAI [2024b] on the dataset to automatically assign initial category tags, which are then relabeled by the expert annotators.
To ensure annotation quality and reliability, we implemented a save-and-replay feature within the tool, enabling annotators to revisit their annotated segments, replay the selected video portion, and iteratively refine or validate their annotations before final submission. We also verify the annotations amongst ourselves. This process ensured that questions accurately aligned with the video context and targeted implicit rather than explicit reasoning. We ourselves have annotated questions with the intent to probe deeper inferential reasoning rather than relying solely on directly observable content. Thus the author annotation process contributed to both the conceptual depth and technical relevance of the dataset. For video selection, we curated a diverse set of 1K creative videos, comprising movies of varied genres and mediums (3D animated and live-action) known for employing narrative techniques such as implied causality, off-screen action, symbolic representation, and indirect storytelling. We prioritized scenes that challenge viewers to make inferences beyond directly visible actions or objects, focusing on content where critical narrative elements are omitted, subtle, or distributed across frames. Statistical characteristics of our collected dataset can be seen in Figure 3.
Table 2: Mean Pairwise Similarity for Question Embeddings.
(lower = more diverse)
Benchmark MPS ($\downarrow$)
MVBench 0.293
TempCompass 0.228
Cinepile 0.216
Movie-QA 0.191
Implicit-QA 0.161
Question Diversity.
To measure question diversity quantitatively, we compute mean pairwise cosine similarity of sentence embeddings between our questions. This metric for text diversity has previously been used in the literature Tevet and Berant [2021]. We utilize the all-MiniLM-L6-v2 model from the Sentence Transformers library to compute text embeddings. The results of this analysis are provided in the Table 2. Lower mean similarity between questions indicates higher diversity, as can be seen our questions have higher diversity compared to prior works. We further provide dataset diversity in terms of genre, movie release timeline, media type in Section C.
@@ 3.2 Dataset Categorization
To ensure comprehensive coverage of the diverse reasoning abilities required for implicit video understanding, we organized the dataset into nine distinct reasoning categories,
@@
each targeting a specific type of implicit inference. Below, we formally define each category of implicit reasoning. We present qualitative examples in Figure 16 to highlight the implicit nature of questions. We also show differences in temporal length between categories in Figure 4, with counting and vertical spatial reasoning videos being the longest.
@@
Viewpoint and Visibility Questions in this category require inferring what is observable from a particular vantage point, whether it be a character’s perspective or camera angle. They must reason about line-of-sight, occlusions, and spatial orientations.
@@
Motivational Reasoning Questions in this category require viewers to infer character motives, or likely future events based on incomplete or indirect visual information. These questions emphasize unstated cause-effect chains within the story.
@@
Figure 6: More Qualitative ImplicitQA examples, targeting distinct implicit‐reasoning dimension.